# FIVB ranking: Misstep in the right direction

### v1

@@ Abstract
This work uses a statistical framework to present and evaluate the ranking algorithm that has been used by Fédération Internationale de Volleyball (FIVB) since 2020. The salient feature of the FIVB ranking is the use of the probabilistic model, which explicitly calculates the probabilities of the games to come. This explicit modeling is new in the context of official ranking, and we study the optimality of its parameters as well as its relationship with the ranking algorithm as such. The analysis is carried out using both analytical and numerical methods. We conclude that, from the modeling perspective, the use of the home-field advantage (HFA)would be beneficial and that the weighting of the game results is counterproductive. Regarding the algorithm itself, we explain the rationale beyond the approximations currently used and explain how to find new parameters which improve the performance. Finally, we propose a new model that drastically simplifies both the implementation and interpretation of the resulting algorithm.
@@ 1 Introduction
The ranking of teams/players is one of the fundamental problems in competitive sports and is used, for example, to declare the champion or to promote and relegate the teams between leagues. In general terms, the ranking is meant to reflect the relative strength of the teams/players, and should have some forecasting capabilities, i.e., to allow one to predict the game results which may be useful for organizing well-balanced tournaments (Csató, 2024).
@@
In this work, we present and evaluate the ranking algorithm, used by the Fédération Internationale de Volleyball (FIVB). Our study is motivated by the modern approach adopted by the FIVB, where six outcomes of the volleyball games have explicit probabilistic models defined in FIVB (2024). To our best knowledge, this is the first officially adopted ranking algorithm to use such an approach and, merely due to this fact, deserves attention in sport analytics. Of course, international volleyball is also a popular sport, and understanding its rating strategy is interesting on its own merit.
@@
The FIVB ranking adopted in 2020 can be classified as power-ranking, where teams are assigned a real-valued parameter called skills (also known as strength or power) and the teams are ranked (ordered) by sorting the skills. This approach departs from the more conventional ranking based on counting of the points associated with the results of the games, and is often considered to be more “fair”.
@@
The power-ranking approach was already adopted by the Fédération Internationale de Football Association (FIFA) to rank the Men’s and Women’s teams, and the analysis made by Szczecinski and Roatis (2022) revealed its strength and showed possible improvements. One of the criticisms of the FIFA ranking made in Szczecinski and Roatis (2022) is that, being based on the Elo rating Elo (2008), it inherits its main drawback, i.e., the lack of an explicit probabilistic model of the game outcomes Szczecinski and Djebbi (2020). From a statistical perspective, this lack of forecasting capability is, indeed, a significant drawback which makes it difficult to evaluate the ranking objectively.
In this regard, the FIVB ranking proposes a radical and modern approach: each of the six outcomes of the volleyball game is assigned a probability that is calculated from the skills (known before the game) using the so-called Cumulative Link (CL) model (Tutz, 2012, Ch. 9.1).
The objective of this work is to present the FIVB ranking in a statistical framework, propose the evaluation methodology, and analyze the optimality of the algorithm’s parameters. Since there are many sports whose games end with multiple outcomes (i.e., more than two), understanding how ranking algorithms may be constructed in such cases will be useful beyond the context of the FIVB ranking.
@@
This work is organized as follows: in Sec. 2, we cast the ranking in the inference context, where the goal is to estimate the skills from the outcomes of the games. This allows us to understand the relationship between the probabilistic model and the ranking algorithm as such. The parameters of the model are assessed in Sec. 3, where, both the analytical and numerical approaches are applied to evaluate the importance of the thresholds in the CL model, the numerical scores used in the FIVB algorithms, the role of the home-field advantage (HFA), and the utility of weights associated with games’ categories. In Sec. 4 we show that, using a logistic function in the CL model (instead of the Gaussian cumulative density function (CDF), currently defined in the FIVB ranking) produces a new, easy to interpret ranking algorithm which, moreover, does not need auxiliary numerical scores which are required in the current FIVB ranking.
@@
We terminate the work in Sec. 6 summarizing the finding and showing the recommended changes to the FIVB ranking algorithm. Overall, we conclude that, from the statistical perspective, using an explicit probabilistic model to build a ranking algorithm is a step in the right direction. On the other hand, we qualify it as misstep because many approximations lead to sub-optimality, which might have been easily avoided.
@@ 2 Model and ranking
Consider the scenario, where, of the $M$ teams, two are selected to face each other in a game. The games are indexed with $t=1,2,\ldots,T$. Teams are indexed with a pair $(i_{t},j_{t})$, where $i_{t},j_{t}\in\{1,\ldots,M\}$. The team $i_{t}$ is called the home-team, and the team $j_{t}$ the away-team. We keep this naming convention even when a game is played on a neutral venue. [^1] A venue is neutral when, during an international tournament, the game is played in the country different from the team $i_{t}$ or $j_{t}$.
The outcome of the $t$-th game $y_{t}\in\mathcal{Y}$ is ordinal in nature, with meaning such as “significance”, which has no numerical value but may be ordered, and, for convenience, we index it with natural numbers $\mathcal{Y}=\{0,\ldots,L-1\}$ in decreasing significance order, i.e., the outcome $y=0$ is the most important and $y=L-1$ is the least important one, where the significance is evaluated from the point of view of the home team $i_{t}$. Quite naturally, the most significant outcome for the home team is the least significant for the away team.
In particular, volleyball games, which we will be interested in, produce $L=6$ possible outcomes “3-0”, “3-1”, “3-2”, “2-3”, “1-3”, “0-3”, where “$k$-$l$” means that the home team won $k$ sets and the away team won $l$ sets; i.e., the first three results mean that the home team won. For the purpose of the derivations, we use indices $y\in\mathcal{Y}$, but it is easier to understand the meaning of the explicit results in the form “$k$-$l$”; i.e., the result $y_{t}=0$ corresponds to the outcome “3-0”, $y_{t}=1$ – to “3-1”, and $y_{t}=5$ – to “0-3”. We will use both and it should not lead to any confusion.
2.1 Rating as statistical inference
The goal of the ranking is to order the teams and, to this end, we assume that they are characterized by intrinsic parameters called “skills” $\theta_{m}\in\mathbb{R},m=1,\ldots,M$. Then, by inferring the skills $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{M}]^{\mathsf{T}}$ from the observed outcomes of the games $\boldsymbol{y}=[y_{1},\ldots,y_{T}]^{\mathsf{T}}$, and ordering them, the ranking is naturally obtained.
@@ 2.1 Rating as statistical inference
$\hat{\boldsymbol{\theta}}$ $=\mathop{\mathrm{argmax}}_{\boldsymbol{\theta}}\log f\Big{(}\boldsymbol{\theta}\big{|}\{y_{t},\boldsymbol{x}_{t}\}_{t=1}^{T}\Big{)}$ (4)
$=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\sum_{t=1}^{T}\ell_{y_{t}}(z_{t})+\rho(\boldsymbol{\theta}),$ (5)
where we use the common assumption that $Y_{t},t=1,\ldots,T$ are independent when conditioned on $\boldsymbol{x}_{t}$ and $\boldsymbol{\theta}$, the negated log-score
@@
is a convenient representation of probability in log domain, and the prior distribution of the skills is defined via $\rho(\boldsymbol{\theta})=-\log f(\boldsymbol{\theta})$.
For example, in the simple case of binary win/loss games $y_{t}\in\{0,1\}$, we may adopt the logistic model,
$\mathsf{P}_{y}(z)$ $=\mathcal{L}\big{(}(-1)^{y}z\big{)},\quad y=0,1$ (7)
$\mathcal{L}(z)$ $=\frac{1}{1+\mathrm{e}^{-z}},$ (8)
and use the Gaussian prior for $\theta_{m}$,
$f(\theta_{m})$ $=\mathcal{N}\Big{(}\theta;0,\frac{1}{\gamma}\Big{)}$ (9)
where
$\mathcal{N}(z;\mu,v)=\frac{1}{\sqrt{2\pi v}}\exp\Big{(}-\frac{1}{2v}(z-\mu)^{2}\Big{)}$ (10)
is the Gaussian probability density function (PDF). Then,
$\rho(\boldsymbol{\theta})$ $=\frac{1}{2}\gamma\|\boldsymbol{\theta}\|^{2}+{\textnormal{Const.}},$. , (11)
and, by finding $\hat{\boldsymbol{\theta}}$ in (4) we implement a logistic regression with $L_{2}$ regularization.
@@
where we use the “loss function” $\ell^{{\textnormal{loss}}}_{y}(z)$ which may, but does not need to, be the same as the log-score $\ell_{y}(z)$ in (6). As we shall see, to simplify calculations, the loss function may be a proxy of the log-score. In this “machine learning” formulation, $\rho(\boldsymbol{\theta})$ is called a regularization function.
Stochastic gradient rating
While (12) provides a conceptual reference framework to analyze the models used for rating and can be applied with a moderate value of $T$ (when it is reasonable to assume that the skills do not change significantly), the online (real-time) rating, where the skills are updated after each game, is more common in practice.
To obtain such a rating algorithm from the previous formulation, it enough to set $\rho(\boldsymbol{\theta})=0$, and solve (12) using the stochastic gradient (SG) algorithm, defined as follows:
@@
When $\ell^{{\textnormal{loss}}}_{y}(z)$ is convex in $z$, for sufficiently large $t$ and appropriately chosen $\mu$, the solution of (15) approximates “well” $\hat{\boldsymbol{\theta}}$. The adaptation step, $\mu$, trades off the convergence speed (which tells how quickly, with $t$, $\hat{\boldsymbol{\theta}}_{t}$ approaches $\hat{\boldsymbol{\theta}}$) against the accuracy (which measures how far $\hat{\boldsymbol{\theta}}_{t}$ is from $\hat{\boldsymbol{\theta}}$). Although this is, admittedly, a vague statement, the precise analysis of the SG solution is not trivial even if some light is shed on this issue, e.g., in Aldous (2017), Jabin and Junca (2015), Szczecinski and Roatis (2022), Gomes de Pinho Zanco et al. (2024). In fact, (Szczecinski and Roatis, 2022, Sec. 3.3) shows that (15) may be seen as an approximation of a nonlinear Kalman filter, which estimates skills at time $t$ from previous observations $y_{1},\ldots,y_{t-1}$.
In this perspective, the skills $\hat{\boldsymbol{\theta}}_{t}$ are approximate solutions to the maximum likelihood (ML) estimation of the skills $\boldsymbol{\theta}$: the approximation is due to the use of the stochastic gradient and due to the use of the loss function, which may approximate the log-score.
@@
Home-field advantage (HFA)and weighting
@@
For example, we may want to take into account the HFA using a binary variable $h_{t}\in\{0,1\}$ that indicates whether the game $t$ is played on the home venue, i.e., in the country of the team $i_{t}$, (then $h_{t}=1$), or is played in the neutral venue (when $h_{t}=0$). The popular model relies on boosting the skills of the home team, or, equivalently, on increasing the values of $z_{t}$ by the HFA parameter $\eta$, i.e.,
@@
where $v_{t}$ is a categorical variable associated with game $t$, $v_{t}\in\{0,\ldots,K-1\}$, and the weight $\xi_{v}\geq 0$ allows us to modulate the relative importance of the term $\ell^{{\textnormal{loss}}}_{y}(z)$: the smaller $\xi_{v_{t}}$ is, the less impact the pair $(z_{t},y_{t})$ will have on the solution $\hat{\boldsymbol{\theta}}$.
In the rating context, $v_{t}$ is called the “prestige” of the game (in the FIVB ranking) or its “importance” (in the FIFA ranking).
@@
Although subjective weighting is not uncommon in statistical literature, see e.g., Hu and Zidek (2001), Ley et al. (2019), it is possible to evaluate the weighting objectively. For example, Szczecinski and Roatis (2022) concluded that the use of weighting is counterproductive in the FIFA ranking.
2.2 FIVB rating algorithm
@@
where $\Phi(z)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{z}\exp(-0.5v^{2})\,\mathrm{d}v$ is the CDF of a zero-mean, unit-variance Gaussian distribution. The model is parameterized with thresholds $\boldsymbol{c}=[c_{-1},c_{0},\ldots,c_{L-1}]$, which are monotonically increasing with $l$, and, to simplify the discussion, may also be assumed symmetric, i.e.,
@@
and always set the first, and the last thresholds as $c_{-1}=-c_{L-1}=-\infty$.
The FIVB ranking defined the thresholds as follows:
$c^{{\textnormal{FIVB}}}_{0}=-c^{{\textnormal{FIVB}}}_{4}=-1.06,$ $c^{{\textnormal{FIVB}}}_{1}=-c^{{\textnormal{FIVB}}}_{3}=-0.394,$ $c^{{\textnormal{FIVB}}}_{3}=0.$ (23)
@@
Figure 1: Functions $\mathsf{P}^{{\textnormal{FIVB}}}_{y}(z)$ in the FIVB rating algorithm, where $y=0$ corresponds to the result “3-0”, $y=1$ to “3-1”, etc.
With the model defined, the FIVB ranking algorithm estimates the skills as follows,(FIVB, 2024):
@@
$3$ $3.50$ Olympic Games qualifying,
@@
Table 2: Numerical values $r^{\textnormal{FIVB}}_{y}$ assigned to the outcomes $y$ in the FIVB rating algorithm.
The above presentation of the FIVB algorithm is meant to be compatible with the formulation in (17) but, of course, the official FIVB presentation in FIVB (2024) uses a different notation, as we explain in Appendix A. For example, instead of defining the step, the scale and the weights, their product $\mu s\xi_{v}$, used in (25) is defined.
The FIVB ranking has an additional rule which imposes a decrease of the rating by 50 points, $\theta_{m,t}\leftarrow\theta_{m,t}-50$, at the end of the year for each team $m$ that did not participate in any competition. The objective is to incentivize teams to play and discourage game avoidance, which might be a strategy used to preserve ranking points. But these are heuristics that cannot be easily put in the statistical framework, and thus we do not model them.

### v2

@@ Abstract
This work presents and evaluates the ranking algorithm that has been used by Fédération Internationale de Volleyball (FIVB) since 2020. The prominent feature of the FIVB ranking is the use of the probabilistic model, which explicitly calculates the probabilities of the future matches results using the estimated teams’ strengths. Such explicit modeling is new in the context of official sport rankings, especially for multi-level outcomes, and we study the optimality of its parameters using both analytical and numerical methods. We conclude that from the modeling perspective, the current thresholds fit well the data but adding the home-field advantage (HFA)would be beneficial. Regarding the algorithm itself, we explain the rationale behind the approximations currently used and show a simple method to find new parameters (numerical score) which improve the performance. We also show that the weighting of the match results is counterproductive.
@@ 1 Introduction
The ranking of teams/players is one of the fundamental problems in competitive sports. It is used, for example, to declare the champion or to promote and relegate teams between leagues, and, in international competitions, to establish the composition of groups e.g., in the qualification rounds of the FIFA World Cup. In other words, the ranking is a tool that allows the governing bodies to manage competitions by “fairly” evaluating the teams. In general terms, the ranking is meant to reflect the relative strength of the teams/players, and may be used for a quick assesment of the competitive landscape and, more consequentially, for tournament design where the strongest teams are scheduled to play in different group phase (Csató, 2024). The formal evaluation of the ranking is based on its ability to predict the results of future matches (Csató, 2024).
@@
In this work, we present and evaluate the ranking algorithm, used by Fédération Internationale de Volleyball (FIVB). Our study is motivated by the modern approach adopted by FIVB, where six-level outcomes of the volleyball matches have explicit probabilistic models defined by FIVB (2024). To our knowledge, among main international sports, [^1] Such as football, cricket, field hockey, tennis, volleybal, table tennis, golf, basketball, ice hockey, or tennis. this is the first officially adopted ranking algorithm to use such an approach and, merely due to this fact, deserves attention in sport analytics. Of course, international volleyball is also a popular sport, and understanding its ranking strategy is interesting on its own merit.
@@
The FIVB ranking adopted in 2020 can be classified as power-ranking, where teams are assigned a real-valued parameter called skills (also known as strength or power) and the teams are ranked (ordered) by sorting the skills. This approach departs from the more conventional ranking based on counting of the points associated with the results of the matches, and is often considered to be more “fair”.
@@
The power-ranking approach was already adopted by Fédération Internationale de Football Association (FIFA) to rank the Men’s and Women’s teams. The analysis of Szczecinski and Roatis (2022) revealed its weaknesses, where one of the criticisms was that the FIFA ranking, being based on the Elo ranking (Elo, 2008), inherits its main drawback, i.e., the lack of an explicit probabilistic model of the match’s outcomes (Szczecinski and Djebbi, 2020). From a statistical perspective, this lack of forecasting capability is, indeed, a significant drawback which makes it difficult to evaluate the ranking objectively. In this regard, the FIVB ranking proposes a radical and modern approach: each of the six outcomes of the volleyball match is assigned a probability that is calculated from the skills (known before the match) using the Cumulative Link (CL) model (Tutz, 2012, Ch. 9.1).
The objective of this work is to reveal the assumptions and simplifications used to derive the FIVB algorithm and to evaluate how well they are applied. In particular, we want to: i) propose the evaluation methodology of the FIVB ranking by casting it in a statistical framework, ii) show approximations used to derive the algorithm, and to iii) assess the optimality of the algorithm’s parameters. In other words, we want to explain how and why the algorithm is working and whether it can be improved in the current general formulation. In this regard, our work follows the approach of Szczecinski and Roatis (2022) which focused on evaluation and understanding of the currently used FIFA ranking. That is, we do not want to propose/analyze new models or algorithms which would be rather in the spirit of many previous works in the area of sport ranking, e.g., in (Karlis and Ntzoufras, 2008), (Egidi et al., 2018), (Ntzoufras et al., 2019), (Gabrio, 2020), (Lasek and Gagolewski, 2021), (Szczecinski, 2022), or (Macrì Demartino et al., 2024).
Since there are many sports with multi-level outcomes (i.e., taking more than two possible values), providing an understanding of how ranking algorithms may be constructed in such cases will be useful beyond the context of the FIVB ranking. We note that the previous works, e.g., (Egidi and Ntzoufras, 2019) or (Ntzoufras et al., 2019), already addressed the issue of modeling in volleyball. In our work, however, we analyze the ranking algorithm currently used by FIVB which, as we show, is not straightforwardly deduced from the model.
@@
This work is organized as follows: in Sec. 2, we cast the ranking in the inference context, where the goal is to estimate the skills from the outcomes of the matches. This allows us to understand the approximate relationship between the probabilistic model and the FIVB ranking algorithm as such. The parameters of the model are assessed in Sec. 3, where, both the analytical and numerical approaches are applied to evaluate the importance of the model thresholds, the numerical scores used in the FIVB algorithms, the role of the home-field advantage (HFA), and the utility of weights associated with matches’ categories. The parameters obtained in Sec. 3, are used in Sec. 4 to assess the performance of the real-time ranking using the FIVB international matches.
@@
We terminate the work in Sec. 5 summarizing the findings and showing the recommended changes to the FIVB ranking algorithm. Overall, we conclude that, from the statistical perspective, using an explicit probabilistic model to build a ranking algorithm is a step in the right direction. On the other hand, we qualify it as misstep because many approximations lead to sub-optimality, which might have been easily avoided.
The repository https://github.com/brbalab/FIVB contains code/data from which the results can be reproduced.
@@ 2 Model and ranking
Consider the scenario, where, of the $M$ teams, two are selected to face each other in a match. The matches are indexed with $t=1,2,\ldots,T$. Teams are indexed with a pair $(i_{t},j_{t})$, where $i_{t},j_{t}\in\{1,\ldots,M\}$. The team $i_{t}$ is called the home-team, and the team $j_{t}$ the away-team. We keep this naming convention even when a match is played on a neutral venue. [^2] A venue is neutral when, during an international tournament, the match is played in the country which is not the one of the team $i_{t}$ or $j_{t}$.
The outcome of the $t$-th match $y_{t}\in\mathcal{Y}$ is ordinal in nature, with meaning such as “importance”, which has no numerical value but may be ordered, and, for convenience, we index it with natural numbers $\mathcal{Y}=\{0,\ldots,L-1\}$ in decreasing order, i.e., the outcome $y=0$ is the most important and $y=L-1$ is the least important one, where the importance is evaluated from the point of view of the home team $i_{t}$. Quite naturally, the most important outcome for the home team is the least important for the away team.
In particular, volleyball matches, which we will be interested in, produce $L=6$ possible outcomes “3-0”, “3-1”, “3-2”, “2-3”, “1-3”, “0-3”, where “$k$-$l$” means that the home team won $k$ sets and the away team won $l$ sets; i.e., the first three results mean that the home team won. For the purpose of the derivations, we use indices $y\in\mathcal{Y}$, but it is easier to understand the meaning of the explicit results in the form “$k$-$l$”; i.e., the result $y_{t}=0$ corresponds to the outcome “3-0”, $y_{t}=1$ – to “3-1”, and $y_{t}=5$ – to “0-3”. We will use both and it should not lead to any confusion.
2.1 Ranking as statistical inference
The goal of the ranking is to order the teams and, to this end, we assume that they are characterized by intrinsic parameters called “skills” $\theta_{m}\in\mathbb{R},m=1,\ldots,M$. Then, by inferring the skills $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{M}]^{\mathsf{T}}$ from the observed outcomes of the matches $\boldsymbol{y}=[y_{1},\ldots,y_{T}]^{\mathsf{T}}$, and ordering them, the ranking is naturally obtained.
@@ 2.1 Ranking as statistical inference
$\hat{\boldsymbol{\theta}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\sum_{t=1}^{T}\ell_{y_{t}}(z_{t})+\rho(\boldsymbol{\theta}),$ (4)
where, the negated log-score
@@
and the prior distribution of the skills is defined via $\rho(\boldsymbol{\theta})=-\log f(\boldsymbol{\theta})$. The common assumption is to use the zero-mean Gaussian model for $\boldsymbol{\theta}$, and then
$\rho(\boldsymbol{\theta})$ $=\frac{1}{2}\gamma\|\boldsymbol{\theta}\|^{2}+{\textnormal{Const.}}$. (6)
@@
where we use the “loss function” $\ell^{{\textnormal{loss}}}_{y}(z)$ which may, but does not need to, be the same as the log-score $\ell_{y}(z)$ in (5). As we shall see, to simplify calculations, the loss function may be a proxy of the log-score. In this “ordinal regression” formulation, $\rho(\boldsymbol{\theta})$ is called a regularization function, and using (6) yields the well-known L2 regularization.
Note that, using $\gamma=0$ and $\ell^{{\textnormal{loss}}}_{y}(z)=\ell_{y}(z)$ we obtain the conventional maximum likelihood (ML) estimation of the skills, which may be appropriate for “sufficiently large” $T$. However, with small/moderate $T$, it is prudent to use regularized form (7). [^3] Regularization is strictly necessary if there are teams whose matches finish with the extreme results “3-0” or “0-3” because then, without regularization, their skills tend to infinity.
Stochastic gradient ranking
We provide (7) to lay out a conceptual reference framework in which we can analyze the models used for ranking. This regression approach can be applied with a moderate value of $T$ (when the skills do not change significantly).
However, as also noted by Csató (2024), practical considerations of simplicity and transparency are more important in the context of sport ranking than the exact formulation. Thus, the online (real-time) ranking, where the skills are updated after each match, is more common in practice and is obtained by solving (7) using the stochastic gradient (SG) algorithm, defined as follows:
@@
When $\ell^{{\textnormal{loss}}}_{y}(z)$ is convex in $z$, for sufficiently large $t$ and appropriately chosen $\mu$, the solution of (10) approximates “well” $\hat{\boldsymbol{\theta}}$. The adaptation step, $\mu$, trades off the convergence speed (which tells how quickly, with $t$, $\hat{\boldsymbol{\theta}}_{t}$ approaches $\hat{\boldsymbol{\theta}}$) against the accuracy (which measures how far $\hat{\boldsymbol{\theta}}_{t}$ is from $\hat{\boldsymbol{\theta}}$). Although this is, admittedly, a vague statement, the precise analysis of the SG solution is not trivial even if some light is shed on this issue, e.g., in (Aldous, 2017), (Jabin and Junca, 2015), (Szczecinski and Roatis, 2022), (Gomes de Pinho Zanco et al., 2024). In fact, (10) may be seen as an approximation of a nonlinear Kalman filter, which estimates skills at time $t$ from previous observations $y_{1},\ldots,y_{t-1}$ (Szczecinski and Roatis, 2022, Sec. 3.3).
In this perspective, the skills $\hat{\boldsymbol{\theta}}_{t}$ are approximate solutions to the ML estimation of the skills $\boldsymbol{\theta}$: the approximation is due to the use of the stochastic gradient and due to the use of the loss function, which may approximate the log-score.
@@
2.2 Integrating exogenous variables
@@
Home-field advantage (HFA)
For example, we may want to take into account the HFA using a binary variable $h_{t}\in\{0,1\}$ that indicates whether the match $t$ is played on the home venue, i.e., in the country of the team $i_{t}$ (then $h_{t}=1$), or is played in the neutral venue (then $h_{t}=0$). The popular model relies on boosting the skills of the home team, or, equivalently, on increasing the values of $z_{t}$ by the HFA parameter $\eta$, i.e.,
@@
where $v_{t}$ is a categorical variable associated with match $t$, $v_{t}\in\{0,\ldots,K-1\}$, and the weight $\xi_{v_{t}}\geq 0$ allows us to modulate the relative importance of the term $\ell^{{\textnormal{loss}}}_{y}(z)$: the smaller $\xi_{v_{t}}$ is, the less impact the pair $(z_{t},y_{t})$ will have on the solution $\hat{\boldsymbol{\theta}}$.
In the ranking context, $v_{t}$ is called the “prestige” of the match (in the FIVB ranking) or its “importance” (in the FIFA ranking).
@@
2.3 Current FIVB ranking algorithm
@@
where $\Phi(z)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{z}\exp(-0.5v^{2})\,\mathrm{d}v$ is the cumulative density function (CDF) of a zero-mean, unit-variance Gaussian distribution.
The model (16) is an extension of the Bradley-Terry model for binary outcomes (Bradley and Terry, 1952) and was also used with ternary outcomes by Glenn and David (1960). The generalization to multi-level outcomes in (16) is known as the ordinal probit (Agresti, 2013, Ch. 8.3); we call it a “CL model” (Tutz, 2012, Ch. 9.1) which is more general allowing us to use non-Gaussian CDF s in the model.
The model is parameterized with thresholds $\boldsymbol{c}=[c_{-1},c_{0},\ldots,c_{L-1}]$, which are monotonically increasing with $l$, and, to simplify the discussion, may also be assumed symmetric, i.e.,
@@
and we always set the first, and the last thresholds as $c_{-1}=-c_{L-1}=-\infty$. Also, for even $L$, we have $c_{\frac{L}{2}-1}\equiv 0$. In particular, in the FIVB model, with $L=6$, we always have $c_{2}\equiv 0$, and $-c_{-1}=c_{5}=\infty$ and thus only two parameters can be set independently: $c_{0}$ and $c_{1}$.
The FIVB ranking defines the thresholds as follows:
$c^{{\textnormal{FIVB}}}_{0}=-c^{{\textnormal{FIVB}}}_{4}=-1.06,$ $c^{{\textnormal{FIVB}}}_{1}=-c^{{\textnormal{FIVB}}}_{3}=-0.394,$ $c^{{\textnormal{FIVB}}}_{2}=0.$ (18)
@@
Figure 1: Functions $\mathsf{P}^{{\textnormal{FIVB}}}_{y}(z)$ in the FIVB ranking algorithm, where $y=0$ corresponds to the result “3-0”, $y=1$ to “3-1”, etc.
With the model defined, the FIVB ranking algorithm (FIVB, 2024) estimates the skills as follows:
@@
$3$ $3.50$ Olympic matches qualifying,
@@
Table 2: Numerical scores $r^{\textnormal{FIVB}}_{y}$ assigned to the outcomes $y$ in the FIVB ranking algorithm.
The FIVB algorithm defined by (20)-(22) is meant to be compatible with the formulation in (12) but, of course, the official FIVB presentation (FIVB, 2024) uses a different notation. For example, instead of defining the step, the scale and the weights, the FIVB ranking defines their product $\mu s\xi_{v}$, which is required in (20); we explain this in detail in Appendix A.
The FIVB ranking has an additional rule, where, at the end of the year, each team $m$ that did not participate in any competition has the value of its skills reduced $\theta_{m,t}\leftarrow\theta_{m,t}-50$. The objective is to discourage match avoidance to preserve the value of the skills (and thus the ranking position). But these are heuristics that cannot be easily put in the statistical framework, and thus we do not model them.